MILLIMETER WAVE SPECTRUM OF THE ^{13}C^{16}O DIMER

L.A. SURIN, D. FOURZIKOV, G. WINNEWISSER, I. Physikalisches Institut, Universität zu Köln, Zülpicher Str. 77, 50937 Köln, Germany; JIAN TANG AND \underline{A.R.W. McKELLAR}, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario K1A 0R6, Canada.

The millimeter wave spectrum of the isotopically substituted CO dimer, (^{13}C^{16}O)_{2} has been studied for the first time, confirming and extending a recent infrared study [1]. A total of 87 transitions in the 77 - 180 GHz region have been assigned and analyzed with a model-independent term value scheme involving 57 rotational levels with J = 0 to 8. The levels can be classified into 7 “stacks” which have symmetry classifications of either A^{+}/B^{−} or A^{−}/B^{+}, and K-values of either 0 or 1. For the normal isotope, symmetry and nuclear spin statistics cause alternate rotational levels to be missing, but for (^{13}C^{16}O)_{2}, all levels are present with an intensity alternation of 1:3 between the A and B symmetries. The four A^{+}/B^{−} stacks have not previously been observed, and the lowest of them establishes the tunneling splitting of (^{13}C^{16}O)_{2} to be 3.769 cm^{−1}, slightly larger than the (^{12}C^{16}O)_{2} value [2] of 3.731 cm^{−1}. Even though a considerable amount of precise experimental data is now available for the CO dimer, we still have little theoretical insight into its structure and tunneling dynamics.

References

[1] A.R.W. McKellar, J. Chem. Phys. 115, 3571 (2001).

[2] J. Tang, A.R.W. McKellar, L.A. Surin, D.N. Fourzikov, B.S. Dumesh, and G. Winnewisser, J. Mol. Spectrosc. 214, 87 (2002).